# Reinforcement Unlearning

### v2

@@ 4.3. Hyperparameter Study
Refer to caption (a) Steps of decremental RL-based Refer to caption (b) Rewards of decremental RL-based Refer to caption (c) Steps of poisoning-based Refer to caption (d) Rewards of poisoning-based
@@
Figure 6. The decremental RL-based and poisoning-based methods in Grid World with different sizes.
@@
In the grid world setting, we extend the size of the environment from $5\times 5$ to $15\times 15$, resulting in a larger grid. Figure 6 visually depicts the impact of this increased environment size on both methods. As illustrated in the figure, we observe that with the expansion of the environment, the discrepancy in rewards between the pre-unlearning and post-unlearning stages is magnified for both methods. The reason is that the larger grid size introduces a greater number of states for the agent to navigate. Thus, unlearning becomes a more challenging task as the agent must modify its learned behavior to adapt to the enlarged environment. Also, this magnification effect can be attributed to the increased number of possible trajectories and interactions in the expanded grid world. After unlearning, the behavior of the agent in the unlearning environment becomes randomized. Consequently, a wider range of possible trajectories and interactions often leads to longer paths taken by the agent, thereby resulting in lower rewards attained. Hence, the discrepancy in rewards between the pre-unlearning and post-unlearning stages becomes more pronounced.
Impact of Poisoning Level.  The hyperparameter, poisoning level, serves as a pivotal factor in evaluating and testing the poisoning-based method exclusively. This parameter governs the quantity of poison introduced to the agent during the unlearning process, enabling us to investigate how the method performs under varying levels of poisoning. Specifically, the poisoning level is measured by the difference between the intended state $s^{\prime}$ and the manipulated state $\hat{s}^{\prime}$. To illustrate, in the grid world context where an agent’s state comprises eight dimensions, a poisoning level of 3 indicates that the two states differ in three dimensions.
Refer to caption (a) Steps of poisoning-based Refer to caption (b) Rewards of poisoning-based
@@
Figure 7. The poisoning-based method in Grid World with different poison levels
@@
Consistent results across the remaining scenarios, illustrated in Figure 8 (aircraft landing), Figure 9 (virtual home), and Figure 10 (maze explorer), further strengthen the observations in the grid world setting. The increased poison during unlearning correlates with improved results. These findings emphasize the significance of the poisoning-based approach in reinforcement unlearning. By strategically introducing poison to prompt the agent to forget specific information, this method offers a promising solution for overcoming challenges in unlearning within complex environments.
Refer to caption (a) Average steps of the poisoning-based method Refer to caption (b) Rewards of the poisoning-based method
Figure 8. The poisoning-based method in Aircraft Landing with different poison levels
Refer to caption (a) Average steps of the poisoning-based method Refer to caption (b) Rewards of the poisoning-based method
Figure 9. The poisoning-based method in Virtual Home with different poison levels
Refer to caption (a) Average steps of the poisoning-based method Refer to caption (b) Rewards of the poisoning-based method
Figure 10. The poisoning-based method in Maze Explorer with different poison levels
4.4. Adaptability Study
Dynamic Environments.  To illustrate the performance of our methods in dynamic environments, where the features and layouts of environments can change during agent training, we introduce a slight modification to the unlearning problem. Specifically, we consider an unlearning environment denoted as $\mathcal{M}_{u}$, and we employ time steps to represent changes in the environment. As time progresses in $t$ steps, the evolution of the unlearning environment can be represented as $\mathcal{M}^{1}_{u},\ldots,\mathcal{M}^{t}_{u}$. Thus, the problem of unlearning $\mathcal{M}_{u}$ transforms into the task of unlearning $\mathcal{M}^{1}_{u},\ldots,\mathcal{M}^{t}_{u}$.
Refer to caption (a) Steps of the four methods Refer to caption (b) Rewards of the four methods
Figure 11. Performance of the four methods in dynamic environments in Grid World
The experimental results, shown in Figure 11, were derived by setting $t=5$ and averaging the outcomes across the five environments. Similar outcomes can be observed for other values of $t$, e.g., $3$ and $8$. The results indicate that even in this dynamic setting, the outcomes of post-unlearning remain notably favorable. The effectiveness is attributed to the carefully designed mechanisms inherent in our reinforcement unlearning methods. Both unlearning methods involve dynamism during their operation. The decremental RL-based method dynamically adjusts the agent’s knowledge, ensuring it remains effective even as the environment undergoes alterations. Similarly, the poisoning-based method introduces dynamism by modifying the unlearning environment, ensuring the agent to perform optimally in the evolving environment.
@@ 4.4. Adaptability Study
Refer to caption (a) Steps of the four methods Refer to caption (b) Rewards of the four methods
Figure 12. Generalization of the four methods in Grid World
@@
Refer to caption (a) Steps of decremental RL-based Refer to caption (b) Rewards of decremental RL-based Refer to caption (c) Steps of poisoning-based Refer to caption (d) Rewards of poisoning-based
@@
Figure 13. The decremental RL-based and poisoning-based methods in Grid World with and without noise
@@
The corresponding results of the grid world setting are presented in Figure 13. Upon analyzing the outcomes, a notable observation emerges: both the decremental reinforcement learning-based and poisoning-based methods exhibit remarkable robustness against external noise. Despite the introduction of noise, the difference in both steps and rewards between the pre-unlearning and post-unlearning states remains nearly unchanged for both methods. This robustness can be attributed to the inherent adaptability and resilience of the unlearning methods. In the case of the decremental RL-based method, the gradual modification of the agent’s policy allows it to withstand minor variations in observations, ensuring that its learned behavior remains stable despite external noise.
Similarly, the poisoning-based method’s strategic use of targeted perturbations enables the agent to develop a more adaptive policy. Thus, the agent’s behavior proves to be less affected by the noise, maintaining its consistency in unlearning undesired knowledge.
We have additionally assessed the unlearning process in two similar environments, delving into the consequences on one environment when unlearning the other. For adherence to page limitations, the detailed results are presented in the appendix.
4.5. Environment Inference Testing
This inference enables us to infer the environment that the agent needs to forget, allowing for a comparison of the inference outcomes before and after unlearning. In Figure 14, we present the results of the aircraft landing setting for the decremental RL-based method.
Figure 14. Inference results in Aircraft
The inference successfully recreates about $50\%$ of the unlearning environment before unlearning. However, after unlearning, this inference result significantly reduces to only $20\%$. This result provides clear evidence of a successful unlearning process. Note that the inference strategy employed in our experiments is not particularly effective. However, its use is solely for the purpose of illustrating the change in the percentage of an environment that can be inferred before and after unlearning. Evaluating our unlearning methods against a more potent inference strategy is left as future research.
@@ 4.5. Environment Inference Testing
The reason behind this success lies in the unlearning methods’ capability to modify the agent’s learned policy effectively. Both of the proposed methods adapt the agent’s behavior to forget specific aspects of the environment while preserving essential knowledge. Thus, the environment inference attack becomes less effective in recreating the forgotten parts after unlearning. Also, the visual comparison highlights the unlearning methods’ efficiency in refining the agent’s policy to eliminate unwanted behaviors. The process of inferring the forgotten environment confirms the success of our unlearning methods in reducing the agent’s reliance on previously learned information and adapting to changes in the environment.
Figure 15. Inference results in Grid World
The results observed in the grid world scenario (Figure 15) are similar to the results in the aircraft landing setting, where the environment inference accuracy is significantly reduced after the unlearning process. In the grid world scenario, as the agent unlearns specific navigation patterns, the environment inference struggles to recreate the forgotten parts accurately. The unlearning process alters the agent’s policy, leading to changes in navigation trajectories. Thus, the inferred environment fails to capture the full complexity and intricacies of the original environment.
Note that these results not only show the accuracy of the inference but also exhibit the associated inference difficulty. A lower inference accuracy indicates a higher level of complexity in deducing information about an environment. It is essential to recognize that a diminished accuracy does not merely signify an error rate; rather, it implies that the inference process demands more time and effort to discern the intricacies of the environment accurately.
Figure 16. The time it took the agent to complete a predefined task before and after unlearning in Virtual Home
In the virtual home setting, where an environment inference is not applicable, we evaluated unlearning performance by measuring the time it took the agent to complete a predefined task before and after unlearning. As illustrated in Figure 16, before unlearning, the agent could swiftly complete the task in $10$ seconds. However, after unlearning, the agent needed $1$ minute and $45$ seconds to accomplish the same task in the same environment, showcasing its diminished performance resulting from the unlearning process.
5. Related Work
@@
Machine Unlearning.  The concept of machine unlearning was initially introduced by Cao et al. (Cao and Yang, 2015). They employed statistical query learning and decomposed the model into a summation form, enabling efficient removal of a sample by subtracting the corresponding summand. Later, Bourtoule et al. (Bourtoule et al., 2021) proposed SISA training, which involves randomly partitioning the training set into multiple shards and training a constituent model for each shard. In the event of an unlearning request, the model provider only needs to retrain the corresponding shard model. Warnecke et al. (Warnecke et al., 2023) shifted the focus of unlearning research from removing samples to removing features and labels. Their approach is based on the concept of influence functions, which allows for estimating the influence of data on learning models. Machine unlearning has also been explored from a theoretical perspective. Ginart et al. (Ginart et al., 2019) introduced the concept of $(\epsilon,\delta)$-approximate unlearning, drawing inspiration from differential privacy (DP) (Dwork, 2006; Ruan et al., 2023; Feng et al., 2023). Subsequently, Guo et al. (Guo et al., 2020) formulated unlearning as certified removal and provided theoretical guarantees. They achieved certified removal by employing convex optimization followed by Gaussian perturbation on the loss function. Gupta et al. (Gupta et al., 2021) considered update sequences based on a function of the published model. They leveraged differential privacy and its connection to max information to develop a data deletion algorithm. Thudi et al. (Thudi et al., 2023) argued that unlearning cannot be proven solely by training the model on the unlearned data. They concluded that unlearning can only be defined at the level of the algorithms used for learning and unlearning.
@@
Reinforcement Learning Security.  While reinforcement unlearning remains an underexplored area, considerable research has been devoted to reinforcement learning security. For instance, studies have extensively investigated the vulnerabilities present in RL systems (Wu et al., 2021) and policy explanation with security applications (Yu et al., 2023). However, it is crucial to note that these studies differ significantly from reinforcement unlearning for three distinct reasons. Firstly, the focus of prior research in reinforcement learning security primarily revolves around identifying and addressing vulnerabilities within the learning process. In contrast, our work centers on the fundamental task of how to effectively forget previously acquired knowledge. Secondly, existing studies in reinforcement learning security commonly aim to train robust agents capable of withstanding diverse adversarial activities. In contrast, our objective is to enable the unlearning of knowledge in a well-trained agent, highlighting a different goal and approach. Lastly, prior research endeavors to explain learned policies, particularly within security applications. In contrast, our research focuses on relearning policies based on revoke requests, giving a novel perspective on policy adaptation.
6. Conclusion

### v5

@@ IV-C Hyperparameter Study
(a) Steps of decremental RL-based
(b) Rewards of decremental RL-based
(c) Steps of poisoning-based
(d) Rewards of poisoning-based
@@
Figure 13: The decremental RL-based and poisoning-based methods in Grid World with different sizes.
@@
In grid world, we extend the size of the environment from $5\times 5$ to $15\times 15$, resulting in a larger grid. Figure 13 visually depicts the impact of this increased environment size on both methods. As illustrated in the figure, we observe that with the expansion of the environment, the discrepancy in rewards between the pre-unlearning and post-unlearning stages is magnified for both methods. The reason is that the larger grid size introduces a greater number of states for the agent to navigate. Thus, unlearning becomes a more challenging task as the agent must modify its learned behavior to adapt to the enlarged environment. Also, this magnification effect can be attributed to the increased number of possible trajectories and interactions in the expanded grid world. After unlearning, the behavior of the agent in the unlearning environment becomes randomized. Thus, a wider range of possible trajectories and interactions often leads to longer paths taken by the agent, thereby resulting in lower rewards attained. Hence, the discrepancy in rewards between the pre-unlearning and post-unlearning stages becomes more pronounced.
Impact of Environment Complexity. The complexity of an environment can be characterized by the presence and arrangement of obstacles within it. By modifying the complexity of the environment, we can assess the adaptability of the proposed methods. In the grid world setting, we examine the impact of increasing the environment complexity by introducing more obstacles, with the environment size maintained at $10\times 10$. Figures 14(a) and 14(b) show the outcomes of the decremental RL-based method.
(a) Steps of decremental RL-based
(b) Rewards of decremental RL-based
(c) Steps of poisoning-based
(d) Rewards of poisoning-based
@@
Figure 14: The decremental RL-based and poisoning-based methods in Grid World with different complexity
However, when employing the poisoning-based method (depicted in Figures 14(c) and 14(d)), an interesting observation emerges. Unlike the decremental RL-based method, the difference in both steps and rewards between the pre-unlearning and post-unlearning stages remains relatively stable even as the number of obstacles increases. The reason behind this intriguing behavior lies in the nature of the poisoning-based approach. When we introduce additional obstacles to the environment, the poisoning-based method operates differently compared to the decremental RL-based approach. Instead of modifying the agent’s learned policy gradually, the poisoning-based method incorporates an element of targeted perturbation. As the number of obstacles increases, the poisoning-based method strategically poisons the agent’s policy by introducing deceptive information during the unlearning process. This targeted perturbation causes the agent’s behavior to deviate from the optimal path more significantly, leading to relatively constant differences in both step count and reward between the pre-unlearning and post-unlearning phases.
Secondly, the poisoning-based method’s targeted perturbations are designed to cause deviations from the optimal path. Thus, the agent’s policy becomes consistently misled in the presence of additional obstacles, leading to a stable performance difference between pre- and post-unlearning.
Impact of Poisoning Level.  The hyperparameter, poisoning level, serves as a pivotal factor in evaluating and testing the poisoning-based method exclusively. This parameter governs the quantity of poison introduced to the agent during the unlearning process, enabling us to investigate how the method performs under varying levels of poisoning. Specifically, the poisoning level is measured by the difference between the intended state $s^{\prime}$ and the manipulated state $\hat{s}^{\prime}$. To illustrate, in the grid world context where an agent’s state comprises eight dimensions, a poisoning level of $3$ indicates that the two states differ in three dimensions. The poisoning level is consistently set to $3$ throughout the experiments, unless otherwise specified.
(a) Steps of poisoning-based
(b) Rewards of poisoning-based
Figure 15: Poisoning-based method in Grid World, different poison levels
@@
Consistent results across the remaining scenarios, shown in Figure 16 (virtual home) and Figure 17 (maze explorer), further strengthen the observations in grid world. The increased poison during unlearning correlates with improved results.
(a) Steps of poisoning-based
(b) Rewards of poisoning-based
Figure 16: Poisoning-based method in Virtual Home, different poison levels
(a) Steps of poisoning-based
(b) Rewards of poisoning-based
Figure 17: Poisoning-based method in Maze Explorer, different poison levels
IV-D Adaptability Study
Dynamic Environments.  To simulate dynamic environments, where the features and layouts of environments can change during agent training, we introduce a slight modification to the unlearning problem. Specifically, considering an unlearning environment $\mathcal{M}_{u}$, we employ time steps to represent changes in the environment. As time progresses in $t$ steps, the evolution of the unlearning environment can be represented as $\mathcal{M}^{1}_{u},\ldots,\mathcal{M}^{t}_{u}$. Thus, the problem of unlearning $\mathcal{M}_{u}$ transforms into the task of unlearning $\mathcal{M}^{1}_{u},\ldots,\mathcal{M}^{t}_{u}$.
(a) Steps of the four methods
(b) Rewards of the four methods
Figure 18: The four methods in dynamic environments in Grid World
The experimental results, shown in Figure 18, were derived by setting $t=5$ and averaging the outcomes across the five environments. Similar outcomes can be observed for other values of $t$, e.g., $3$ and $8$. The results indicate that even in this dynamic setting, the outcomes of post-unlearning remain notably favorable. The effectiveness is attributed to the carefully designed mechanisms inherent in our reinforcement unlearning methods. The decremental RL-based method dynamically adjusts the agent’s knowledge, ensuring it remains effective even as the environment undergoes alterations. Similarly, the poisoning-based method introduces dynamism by modifying the unlearning environment, ensuring the agent to perform optimally in the evolving environment.
@@ IV-D Adaptability Study
(a) Steps of the four methods
(b) Rewards of the four methods
Figure 19: Generalization of the four methods in Grid World.
@@
(a) Steps of decre. RL-based with noise
(b) Steps of decre. RL-based without noise
(c) Rewards of decre. RL-based with noise
(d) Rewards of decre. RL-based without noise
(e) Steps of poisoning-based with noise
(f) Steps of poisoning-based without noise
(g) Rewards of poisoning-based with noise
(h) Rewards of poisoning-based without noise
@@
Figure 20: The decremental RL-based and poisoning-based methods in Grid World with and without noise
@@
The corresponding results of the grid world setting are presented in Figure 20. Upon analyzing the outcomes, a notable observation emerges: both the decremental RL-based and poisoning-based methods exhibit remarkable robustness against external noise. Despite the introduction of noise, the agent’s performance in both the unlearning and remaining environments remains consistently stable across both methods, showing no significant difference compared to conditions without noise. This robustness can be attributed to the inherent adaptability and resilience of the unlearning methods. In the decremental RL-based method, the gradual modification of the agent’s policy allows it to withstand minor variations in observations, ensuring that its learned behavior remains stable despite external noise. Similarly, the poisoning-based method’s strategic use of targeted perturbations enables the agent to develop a more adaptive policy. Thus, the agent’s behavior proves to be less affected by the noise, maintaining its consistency in unlearning undesired knowledge.
The results observed in the virtual home (Figure 21) and maze explorer (Figure 22) scenarios closely align with those in the grid world setting, showcasing the robustness of both the decremental RL-based and poisoning-based methods against external noise. In these two scenarios, the agent’s behavior continues to exhibit consistent patterns even when external noise is introduced during both the training and unlearning process. This consistency is crucial for practical real-world applications, where agents must maintain their adaptability and performance despite uncertainties and disturbances.
(a) Average steps of the decremental RL-based method with noise
(b) Average steps of the decremental RL-based method without noise
(c) Rewards of the decremental RL-based method with noise
(d) Rewards of the decremental RL-based method without noise
(e) Average steps of the poisoning-based method with noise
(f) Average steps of the poisoning-based method without noise
(g) Rewards of the poisoning-based method with noise
(h) Rewards of the poisoning-based method without noise
Figure 21: The decremental RL-based and poisoning-based methods in Virtual Home with and without noise
(a) Average steps of the decremental RL-based method with noise
(b) Average steps of the decremental RL-based method without noise
(c) Rewards of the decremental RL-based method with noise
(d) Rewards of the decremental RL-based method without noise
(e) Average steps of the poisoning-based method with noise
(f) Average steps of the poisoning-based method without noise
(g) Rewards of the poisoning-based method with noise
(h) Rewards of the poisoning-based method without noise
Figure 22: The decremental RL-based and poisoning-based methods in Maze Explorer with and without noise
Computational Adaptability. To assess the computational adaptability of the four methods, we initially examined their computational overhead in scenarios where one environment is unlearned from a set of twenty. This evaluation was then expanded to the scenario of unlearning ten environments from a hundred. The outcomes are detailed in Table II.
TABLE II: Computation Overhead of the Four Methods in Grid World (seconds)
Methods Computation time of unlearning 1 environment Computation time of unlearning 10 environments
Decremental RL $18.80s$ $64.83s$
Poisoning $20.17s$ $73.56s$
LFS $198.62s$ $746.25s$
Non-transfer LFS $196.97s$ $742.51s$
In comparison with the two baseline methods, the proposed decremental-RL and poisoning approaches show greater computational efficiency due to their less exhaustive exploration of the environments. Specifically, the decremental RL-based method is more efficient than the poisoning-based method. For instance, when the number of unlearning environments increases to $10$, the poisoning-based method requires $73.56$ seconds, whereas the decremental RL-based method only takes $64.83$ seconds, approximately $12\%$ faster. This efficiency stems from the decremental RL-based method’s straightforward policy adjustments impacting rewards directly without the extensive need to recompute environmental dynamics, unlike the poisoning-based method, which involves computationally intensive alterations to the transition dynamics.
IV-E Privacy Study
The privacy study is conducted using recommendation systems, where the key indicator is recommendation accuracy. Higher accuracy indicates a deep understanding of users’ preferences and habits, reflecting potential privacy risks. Therefore, after a user revokes their data, the system’s unlearning effectiveness is demonstrated by a decrease in recommendation accuracy for that user, suggesting the recommender has effectively forgotten the user’s preferences and is now generating recommendations randomly.
TABLE III: Comparison of Recommendation Accuracy and Rewards Before and After Unlearning
Methods Performance for the unlearned user Average performance for the remaining users
Accuracy Reward Accuracy Reward
Before Unlearning $92.07\%$ $41.42$ $91.52\%$ $39.9$
Decremental RL $68.63\%$ $20.03$ $90.89\%$ $38.82$
Poisoning $64.41\%$ $18.25$ $91.43\%$ $37.17$
Table III presents the recommendation results before and after unlearning. Initially, the recommendation accuracy for the unlearned user is high at $92.07\%$, indicating a strong understanding of the user’s preferences. However, after unlearning, accuracy significantly decreases to $68.63\%$ and $64.41\%$ for the decremental RL-based and poisoning-based approaches respectively, suggesting effective unlearning and enhanced user privacy protection. This shift towards randomness in recommendations is further evidenced by the dramatic drop in rewards from $41.42$ to $20.03$ and $18.25$, indicating dissatisfaction with the recommendations after unlearning. Importantly, recommendations to other users remain consistent, highlighting the targeted nature of the proposed unlearning methods.
IV-F Safety Study
Since conventional machine unlearning can introduce security issues, such as reducing the adversarial robustness of unlearned models [49], we aim to investigate whether reinforcement unlearning might lead to similar vulnerabilities. Our study focuses on a safety-critical scenario: aircraft landing. In this scenario, we measure safety performance by the number of collisions the agent encounters with obstacles. An increase in collisions after unlearning would indicate a higher level of safety concerns, suggesting that reinforcement unlearning could potentially compromise the safety of the system in ways analogous to the security issues observed in conventional machine unlearning.
(a) The number of collisions under the decremental RL-based method
(b) The number of collisions under the poisoning-based method
Figure 23: Collision count before and after unlearning in Aircraft Landing.
The results shown in Figure 23 reveal that after unlearning, the number of collisions under both methods has tripled compared to before unlearning. This substantial increase highlights the safety risks introduced when the agent forgets critical features of the environment. In safety-critical scenarios like aircraft landing, this observation underscores the need for extreme caution when implementing unlearning strategies. The trade-off between ensuring privacy and maintaining safety must be judiciously managed, as compromising safety can have serious consequences. Therefore, in contexts where safety is paramount, the approach to unlearning must prioritize minimal impact on operational safety.
IV-G Environment Inference Testing
This inference enables us to infer the environment that the agent needs to forget, allowing for a comparison of the inference outcomes before and after unlearning. We incorporate the $l_{0}$ distance as a quantitative measure to assess the inference results. Specifically, consider the unlearning environment as $\mathcal{M}_{u}$, and the inferred environments before and after unlearning as $\mathcal{M}_{u}^{before}$ and $\mathcal{M}_{u}^{after}$, respectively. The $l_{0}$ distance between two environments is defined as the number of differing dimensions between them.
Figure 24: Inference results in Grid World
In Figure 24, we present the results of the grid world setting for the decremental RL-based method. The inference successfully recreates $50\%$ of the unlearning environment before unlearning, resulting in $l_{0}(\mathcal{M}_{u},\mathcal{M}_{u}^{before})=12$ for this $5\times 5$ environment. However, after unlearning, this inference result significantly reduces to only $20\%$, yielding $l_{0}(\mathcal{M}_{u},\mathcal{M}_{u}^{before})=20$. This result provides clear evidence of a successful unlearning process. Note that the inference results before unlearning may not appear fully accurate due to the specific inference approach used, which is only intended to verify the effectiveness of the unlearning process, rather than to demonstrate an adversary’s capabilities. The primary objective here is to assess how well the unlearning method obscures the agent’s prior knowledge of the environment.
@@ IV-G Environment Inference Testing
The reason behind this success lies in the unlearning methods’ capability to modify the agent’s learned policy effectively. Both of the proposed methods adapt the agent’s behavior to forget specific aspects of the environment while preserving essential knowledge. Thus, the environment inference becomes less effective in recreating the forgotten parts after unlearning. Also, the visual comparison highlights the unlearning methods’ efficiency in refining the agent’s policy to eliminate unwanted behaviors. The process of inferring the forgotten environment confirms the success of our unlearning methods in reducing the agent’s reliance on previously learned information and adapting to changes in the environment.
V Threats to Validity
Definition Validity. The study defines unlearning in a specific way (minimizing performance in the unlearning environment). Alternative definitions of unlearning might yield different results. Moreover, the defined reward function might not capture all aspects of the agent’s behavior and performance, leading to incomplete assessments of unlearning efficacy.
Methodology Validity. The effectiveness of the proposed unlearning methods may be dependent on the specific RL algorithms used in the study. Additionally, the unlearning process might be influenced by factors not accounted for in the study. For example, in recommendation systems, user preferences can change over time, making it challenging for the agent to provide accurate recommendations. While we have tested various RL algorithms and incorporated environmental dynamism into our experiments, our evaluations are not exhaustive in capturing all potential variables.
Generalization Validity. The experiments were conducted in simulated environments, which may not fully represent the complexities of real-world scenarios. Furthermore, the metrics used to evaluate unlearning performance might not capture all aspects of the agent’s behavior in real-life applications.
VI Related Work
@@
Machine Unlearning.  The concept of machine unlearning was initially introduced in [3]. They employed statistical query learning and decomposed the model into a summation form, enabling efficient removal of a sample by subtracting the corresponding summand. Later, Bourtoule et al. [2] proposed SISA training, which involves randomly partitioning the training set into multiple shards and training a constituent model for each shard. In the event of an unlearning request, the model provider only needs to retrain the corresponding shard model. Warnecke et al. [43] shifted the focus of unlearning research from removing samples to removing features and labels. Their approach is based on the concept of influence functions, which allows for estimating the influence of data on learning models. Machine unlearning has also been explored from a theoretical perspective. Ginart et al. [12] introduced the concept of $(\epsilon,\delta)$-approximate unlearning, drawing inspiration from differential privacy (DP) [6, 31, 9]. Then, Guo et al. [13] formulated unlearning as certified removal and provided theoretical guarantees. They achieved certified removal by employing convex optimization followed by Gaussian perturbation on the loss function. Gupta et al. [14] considered update sequences based on a function of the published model. They leveraged differential privacy and its connection to max information to develop a data deletion algorithm. Thudi et al. [37] argued that unlearning cannot be proven solely by training the model on the unlearned data. Instead, unlearning can only be defined at the level of the algorithms used for learning and unlearning.
@@
Reinforcement Learning Security.  While reinforcement unlearning remains an underexplored area, considerable research has been devoted to reinforcement learning security [44, 47]. For instance, Chen et al. [5] proposed a backdoor detection and removal approach for DRL via unlearning, involving the re-initialization of the top $L$ neurons’ weights to erase the most vulnerable shortcuts. However, these studies differ significantly from reinforcement unlearning for three reasons. Firstly, the focus of these studies revolves around identifying and addressing vulnerabilities within the learning process. In contrast, our work centers on the task of effectively forgetting previously acquired knowledge. Secondly, these studies commonly aim to train robust agents capable of withstanding diverse adversarial activities. Our goal, however, is to develop techniques for unlearning knowledge in RL agents, allowing them to adapt their policies based on unlearning requests. Lastly, these studies endeavor to explain learned policies, particularly within security applications. In contrast, our research extends beyond security concerns to address broader issues related to policy adaptation and forgetting in reinforcement learning.
VII Conclusion